OpenStudy (the_fizicx99):

My LaTeX Thingy - Most likely a fail :3

OpenStudy (the_fizicx99):

\(Hello!\)

OpenStudy (the_fizicx99):

\(\cal\color{blue}{Ello ~ There!!}\)

OpenStudy (the_fizicx99):

\(\dfrac{1}{2}\) + \(\dfrac{59}{23}\) = \(\dfrac{23}{46}\) + \(\dfrac{118}{46}\) = \(\dfrac{141}{46}\) = and I'm stuck can't get the 3 3/46 \(\tt(D;)\)

OpenStudy (solomonzelman):

What? for real stock with that if you know how to write a fraction? 3\frac{3}{46} there is the latex, and it looks like \[3 \frac{3}{46}\] if you are typing words in side equation editor or other latex, use ~ as a space, the more of ~ you use, the greater the space will be.

OpenStudy (solomonzelman):

Also \[\huge\color{blue}{ 3 \frac{1}{2} }\] use, \huge\color{blue}{ 3 \frac{1}{2} }

OpenStudy (solomonzelman):

you can change color from blue, to red green purple black and so forth.

OpenStudy (the_fizicx99):

So, \(\LARGE\color{purple}{ 3 \frac{3}{46} }\)

OpenStudy (solomonzelman):

Yup!

OpenStudy (the_fizicx99):

How would I write exponents

OpenStudy (solomonzelman):

^{ / } ^ { \frac{x}{2} }

OpenStudy (solomonzelman):

you can paste this into the huge latex too.

OpenStudy (the_fizicx99):

I meant like 2^2

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\[\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid purple ]{ 5^ \frac{1}{3} }}}\] \huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid purple ]{ 5^ \frac{1}{3} }}}

OpenStudy (solomonzelman):

this is also a different way, with a box, you can (carefully) try switching colors.

OpenStudy (solomonzelman):

If you want more, ask.....

OpenStudy (the_fizicx99):

Oh sorry I was away, thanks !!

OpenStudy (solomonzelman):

It's fine, yw!

OpenStudy (the_fizicx99):

:)

OpenStudy (solomonzelman):

\(\Huge{\color{blue}{\bigstar!!}}\)

OpenStudy (solomonzelman):

Maybe you can help me with my latex question below?

OpenStudy (anonymous):

blue\color{#FF0000}█